OpenStudy (anonymous):

help Liam deposited $1,859.69 in a savings account that earns 2.4% simple interest. What will Liam’s account balance be in 9 months?

OpenStudy (anonymous):

what formula should I use A=P(1+r/n)^nt

OpenStudy (anonymous):

"....2.4% simple interest " ==> I = P r t

OpenStudy (anonymous):

so 1859.69(0.024)(9)

OpenStudy (anonymous):

1859.69(0.216)

OpenStudy (anonymous):

is 401.69304 and rounded would be 401.70

OpenStudy (anonymous):

t = .75

OpenStudy (anonymous):

1859.69(0.024)(.75) so 33.47442

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

ok what am I doing wrong?

OpenStudy (anonymous):

33.47442 ( correct)

OpenStudy (anonymous):

No I am doing something wrong because my answer choices are 1,859.69 $1,893.16 $1,904.32 $2,261.38

OpenStudy (anonymous):

Maybe I am using wrong formula

OpenStudy (anonymous):

B = $1,859.69 + 33.47442 = .....

OpenStudy (anonymous):

1893.16

OpenStudy (anonymous):

all you do is key into the calculator

OpenStudy (anonymous):

thanks. This is all confusing to me. I have examples in my note book but whenI follow them I get something totally different

OpenStudy (anonymous):

I'm glad it works out :)

OpenStudy (anonymous):

2.4% is yearly, so what you wanna do is 1859.69 + 1859.69(0.024/12)(9) = 1893.16